OpenStudy (realitywillslapyou):

Is this correct? 12 + 4x ≤ 5x -2 12 -x ≤ -2 -x ≤ -14

OpenStudy (jamesj):

Yes, but we would never express the answer that way. You want an inequality involving only x, not -x, hence .... x ....

OpenStudy (realitywillslapyou):

Oh, okay so it's x ≤ -14?

OpenStudy (jamesj):

No. Careful. a < b <=> -a > -b

OpenStudy (realitywillslapyou):

?

OpenStudy (jamesj):

if \[ -x \leq -14 \] that can NOT mean that \( x \leq -14 \). For example, when x = 15, \( -x = 15 \leq -14 \) is true, but \( x =15 \leq -14 \) is definitely false.

OpenStudy (realitywillslapyou):

Okay, so is all of it wrong?

OpenStudy (jamesj):

No. What I am saying is you want to have an inequality involving only x, not -x. The inequality \[ -x \leq -14 \] is equivalent to \[ x \geq 14 \]

OpenStudy (realitywillslapyou):

Oh, okay I get it now. So I should shade it to the left, correct?

OpenStudy (jamesj):

No. x is greater or equal to 14, so it is that part of the number line to the right of 14, including 14 itself.

OpenStudy (realitywillslapyou):

Closed circle shade to the right?

OpenStudy (jamesj):

Yes, as the shaded circle indicates you include that value, in this case 14.

OpenStudy (jamesj):

btw, whenever you do these things you should double check with the original expression. Put in x = 14 and x = 15 and make sure you get the results you expect.

OpenStudy (realitywillslapyou):

Okay, thank you so much! o( ̄∇ ̄o)